OpenStudy (anonymous):

Use the associative law of addition to write an equivalent expression. (m+np)+r

OpenStudy (unklerhaukus):

Associativity \[(a+b)+c=a+(b+c)\]

OpenStudy (anonymous):

i dont get it

OpenStudy (cwrw238):

- what it means is whichever 2 terms you put in the parentheses the result is the same so for ( m + np) + r: (m + np) + r = m + (np + r)

OpenStudy (anonymous):

thank you